Is (√2, 1, 1/2) ε Q×Z×R
1
Expert's answer
2021-08-29T16:43:06-0400

Since "\sqrt 2 \notin Q" then "\left( {\sqrt 2 ;\,1;\,\frac{1}{2}} \right) \notin Q \times Z \times R"

Answer: "\left( {\sqrt 2 ;\,1;\,\frac{1}{2}} \right) \notin Q \times Z \times R"